THE NATURE OF ACTINYL-LIGAND BINDING: AnX_{2}^{2+} IONS (X= S, Se, Te) VALENCE-ISOELECTRONIC WITH AnO_{2}^{2+}

\underline{N. ISMAIL}, C. J. MARSDEN, J.-L. HEULLY and T. SAUE , IRSAMC, Université P. Sabatier, Toulouse, France.

There is substancial current interest in theoretical studies of the actinides, which represent one of the main current challenges for computational chemistry. We have shown that 1-component pseudopotential methods are satisfactory for the uranyl ion; in particular, DFT methods are highly promising. There does not appear to have been any previous work, either experimental or theoretical, on AnX_{2}^{2+} ions containing the heavier Group 16 elements (S, Se, Te). We find that their electronic and geometrical structures bear little resemblance to those of AnX_{2}^{2+}. They have quintet ground states for An=U, and are strongly bent, with X-U-X angles of around 50 degres and substantial X-X bonding. As a result of the differences beetween size and energy of the valence orbitals of S compared to those of O, the predominant bonding interactions in US_{2}^{2+} involve the 6d orbitals on U rather than 5f. While the actinyls are all geometrically similar, the shapes of US_{2}^{2+} and P uS_{2}^{2+} differ appreciably. This observation, together with the novel bonding involved, suggests that these species may have applications as separation agents for the actinides.